# AI4Research: A Survey of Artificial Intelligence for Scientific Research

### v1

@@ 5.4.2 Pre-Experiment Estimation
predicts quantitative values or trends of experimental outcomes, such as estimating drug concentration effects, determining whether a compound affects cellular activity, and assessing protocol feasibility [151]. (1) Deep-Learning Methods: With the rise of deep learning, hierarchical prediction models have emerged [833]. Li et al. [437] incorporated physical equations to predict pharmacokinetic parameters, reducing data requirements and enhancing noise robustness. More recently, Li et al. [438] proposed a dual-matching framework, combining hierarchical molecular alignment with meta-learning, which showed significant improvements in drug feature estimation. (2) LLM-Augmented Methods: More recently, with the advent of LLM capabilities, Zhang et al. [901] demonstrate successful LLM-assisted evaluative prediction, a method that has been further extended in subsequent studies. Notably, Luo et al. [515] integrated BrainGPT into neuroscience literature retrieval, outperforming domain experts in evaluating experimental esitmation. Wen et al. [817] developed a system combining fine-tuned GPT-4.1 with a paper retrieval agent, which outperformed 25 human experts in evaluating experimental predictions.
@@ 5.4.3 Experiment Management
5.4.4 Experimental Conduction
Experimental conduction refers to the application of AI techniques in executing and managing scientific experiments. This process is essential for automating workflows, ensuring experiments are carried out efficiently and accurately. The primary aim of experimental conduction is to reduce human involvement in the experimental process. It can be further divided into two categories:
@@ 5.4.4 Experimental Conduction
uses AI to streamline the design, training, and evaluation of ML models, reducing dependence on human expertise by covering the entire pipeline from preprocessing to hyperparameter optimization [897, 758, 253, 911, 574, 927]. Typically, Wang et al. [799] present a community-driven sandbox allowing agents to write code, browse the web, and coordinate through an event-stream API. For Kaggle challenges, Li et al. [457] propose an iterative, collaborative multi-agent system that incorporates debugging and unit testing across the competition pipeline. AIDE [174] utilizes a tree-search loop to generate, evaluate, and refine solutions, achieving a bronze medal in Kaggle competitions. At the research level, Li et al. [441] formalize a three-phase LLM agent workflow, idea generation, implementation, and execution, to automate experiments. Zhao et al. [921] and Liu et al. [481] present a multi-agent method for extracting model variables from scientific texts, significantly improving experimental reproduction accuracy.
@@
Recent advancements in the planning and reasoning capabilities of LLMs have led to their use in simulating experimental results [357, 882, 414] and even direct conduct real-world experiments [122, 656, 200]. Real-world experimental simulation & conduction generally employ four strategies: (1) Self-Improvement: Models iteratively refine their performance based on feedback [311, 680, 908, 593, 877]. For example, Siddiqui et al. [685] enhance functional approximation through iterative knowledge application. Further refinement occurs through analytical insights [441, 734, 36] and hyperparameter tuning [566, 492, 893]. (2) Multi-Agent Interaction: Models simulate collaborative research teams by assigning roles such as experimenter, analyst, or critic [36, 231, 710, 305, 699]. For instance, MechAgent [565], Researchcodeagent [216] and The AI Scientist [507, 848] automate experiments through multi-agent collaboration, with LLMs acting as proxies in fields like computer science [507, 848], social science [550, 530], and physical science [898]. (3) External Tool Integration: Researchers enhance model capabilities by linking them to databases, APIs, and other tools during experiments [276, 618, 663]. For example, Boiko et al. [68] integrate internet search, code execution, and automation into a GPT-4 system. Studies like ChemCrow [518] and Crispr-GPT [314] support chemistry and gene editing experiments through massive specialized toolchains. (4) Specific Fine-Tuning: A growing body of work explores the fine-tuning of specific models to improve experimental simulations. For instance, Cui et al. [150] present a transformer-based model trained on large single-cell transcriptomic datasets, achieving state-of-the-art accuracy in cell-type annotation and in silico perturbation response [489].
@@ 6 AI for Academic Writing
AI for Academic writing involves the use of AI techniques to assist researchers or generate from scratch in drafting, editing, and formatting scientific manuscripts [371]. With the development of deeper interaction between human and LLMs, human and LLMs are quickly shaping each other’s better writing habits [227, 86, 933]. As shown in Figure 6, it contains two main categories: Semi-Automatic Academic Writing (§$\S$ 6.1) and Full-Automatic Academic Writing (§$\S$ 6.2).
@@ 6.1.2 Assistance During Manuscript Writing
serves as an effective means of conveying experimental data and analysis [55]. Recent advancements in AI research on automatic scientific figure generation have led to significant breakthroughs [894, 293, 747, 635]. Rodriguez et al. [643] introduce the FigGen model, which maps textual descriptions to complex academic figures with high fidelity. Beyond direct image generation, several studies have explored programming scientific diagrams using Python [920, 913, 131, 551], SVG [645], or tikz [55, 54] for better figure quality. However, figures alone cannot fully convey results; captions are essential for ensuring that readers understand each figure [90, 507, 867]. To address this, Hsu et al. [300] develop SciCapenter, an interactive system that generates multiple caption candidates, scores them, and quality-checks each, assisting authors in selecting the most optimal phrasing. MLBCAP [375] utilize multiple LLMs to collaborate on chart-based title generation. Additionally, frameworks like AI Scientist [507, 848] can identify key data from experimental logs, generate figures with captions, and integrate them into authoring tools, advancing end-to-end AI-driven scientific visualization.
@@ 7 AI for Academic Peer Reviewing
Peer reviewing plays a crucial role in enhancing the quality of academic papers. However, it is often hindered by delays, time demands, and growing academic workloads [469, 384, 753, 946]. To address these challenges and improve dissertation quality, researchers are exploring the integration of AI into the review process [880, 490, 573, 399, 523]. As shown in Figure 7, it contains three main categories: Pre-Review (§$\S$ 7.1), In-Review (§$\S$ 7.2) and Post-Review (§$\S$ 7.3).
@@ 7.1.1 Desk-Review
As manuscript submissions to academic journals increase, editorial offices face heavier workloads during the desk-review stage. To address this, major publishers have introduced AI-driven tools, such as automated keyword extraction, topic matching, and preliminary scoring, to enhance efficiency, shorten turnaround times, and reduce manual screening [176, 426, 195]. For example, Elsevier’s Evise and Editorial Manager (EM) systems use indexing and extract terms to route manuscripts to the appropriate subject areas and editorial teams [752]. Similarly, IEEE’s Manuscript Central (built on ScholarOne) combines metadata, author-provided keywords, and an academic-network reviewer discovery tool for more accurate matching [698]. Springer’s SNAPP system and Nature’s AI-assisted triage tool also demonstrate AI’s impact on desk-review workflows [560]. Additionally, Díaz et al. [169] develop AnnotateGPT, which generates annotations to help editors quickly assess a manuscript’s scope and quality, further speeding up the review process.
@@ 7.2 In-Review
Table 5: Comparison of expert and LLM review performance based on Shin et al. [679], where “GPT-4o-Finetuned” refers to GPT-4o finetuned with review data using the finetune-API. KL divergences are calculated from the average of four focus distributions (strength/target, weakness/target, strength/aspect, weakness/aspect) between expert and LLM reviews. F1 scores for overall performance, strength, and weakness are derived by comparing the (target, aspect) sets between expert and LLM reviews. Text similarity metrics are computed to assess the alignment between LLM and expert reviews. Results are sourced from Shin et al. [679]. The bolded contents indicate the highest performance for each metric.
7.2.1 Peer-Review
@@
integrates textual comments and numeric scores into a single review output that mirrors real-world peer review workflows [678, 415]. There are three main paradigms for optimizing unified peer-review generation: (1) Single-Agent Optimization: A straightforward approach is to optimize a single agent through deeper analysis [358, 940, 327]. Shin et al. [679] observe that, by comparing the focus distributions of LLMs and human experts, off-the-shelf LLMs tend to prioritize technical validity in paper reviews while underemphasizing novelty. To address this, Tyser et al. [762] enhance the review system with a suite of review documents to reduce risks of misuse, score inflation, overconfident assessments, and uneven distributions. Additionally, Zhu et al. [940], Zhang and Abernethy [899] incorporate deeper reasoning via reasoning LLMs to improve review quality. (2) Iterative Refinement Optimization: High-quality feedback is often ensured through hierarchical quality control and multi-round refinement loops [59]. Wu et al. [828] propose an LLM-driven pipeline with hierarchical verification, producing literature surveys that match human-authored reviews. Kirtani et al. [380] introduce standardized evaluation metrics and a self-refinement cycle to align LLM-generated reviews with human accuracy and analytical depth. (3) Multi-Agent Optimization: To further enhance feedback reliability, some studies adopt multi-agent frameworks [739, 321, 220, 570]. D’Arcy et al. [157] divide manuscripts into modules for specialized agents, leading to higher-quality feedback than single-agent systems. CycleResearcher [820] and TreeReview [97] apply reinforcement learning to simulate iterative review rounds and structured agent interactions, enhancing collaboration. Furthermore, Taechoyotin and Acuna [735] propose multi-objective reinforcement learning to optimize unified peer review, while Taechoyotin et al. [736] extend multi-agent scientific reviews to multimodal scenarios.
7.2.2 Meta-Review
@@ 7.3.1 Influence Analysis
Influence analysis seeks to predict the future scholarly impact of a paper, most commonly measured by citation count, by evaluating its intrinsic characteristics [694, 325]. Early approaches predominantly rely on external metadata or handcrafted features, such as author reputation and journal impact factor [172, 737, 944]. In contrast, recent methods leveraging LLMs offer the advantage of directly inferring a work’s innovativeness from its narrative. For instance, Zhao et al. [918] frame influence prediction as a regression task, fine-tuning an LLM on titles and abstracts to generate a time- and field-normalized impact score, effectively addressing the cold-start problem. Similarly, the HLM-Cite framework [275] adopts a two-stage approach: first, an embedding model retrieves a set of candidate citations from a large corpus, followed by a generative LLM that performs fine-grained reasoning and re-ranking to identify the most relevant references. Empirical studies [515] suggest that these content-based methods can even surpass human experts in predicting research outcomes in fields such as neuroscience.
@@ 8 Application of AI for Research
As shown in Figure 8, it contains three main categories: AI for Natural Science Research (§$\S$ 8.1), AI for Applied Science and Engineering Research (§$\S$ 8.2) and AI for Social Science Research (§$\S$ 8.3).
@@ 8.1.3 AI for Chemistry & Materials Research
seeks to identify optimal or novel material compositions in virtual or automated experimental setups while minimizing the number of required experiments [46, 630, 669]. Specifically, Chen et al. [101] introduce MEGNet, demonstrating that graph neural networks can achieve density functional theory-level accuracy for both molecular and crystalline properties. Li et al. [450] employ two-stage Bayesian optimization to screen 560 organic photocatalysts using only 2.4% of the experimental conditions, thereby obtaining significantly improved performance. Ekosso et al. [186] combine low-cost robotic platforms with high-throughput microscopy and Gaussian process models to map vesicle formation processes. More recently, Szymanski et al. [734] implement a robot-machine-learning platform that accelerate compound discovery and identify 41 novel inorganic materials within 17 days.
@@ 8.3.2 AI for Psychology Research
Research methodology focuses on the design, implementation, and validation of psychological experiments to ensure validity and reproducibility [757, 427].
@@ 9.1.2 Table & Chart Scientific Comprehension
In the domain of reasoning based on charts and tables, a number of benchmarks have emerged to evaluate the ability of LLMs in both structural and logical comprehension. Early works, such as ChartQA [535], CharXiv [812], ChartX [837], and NovaChart [301], focus on assessing LLMs’ performance in answering questions related to charts, utilizing both synthetic and real-world data. On the other hand, benchmarks like SUC [718], TableBench [832], and ToRR [28] emphasize the evaluation of LLMs’ structural understanding, rather than content comprehension, across various tasks such as form interpretation, numerical reasoning, and textual analysis.
@@ 9.2 AI for Academic Survey
In the task of generating academic surveys, several representative public corpora stand out due to their distinct characteristics in terms of scale, domain, and structure. To evaluate the capabilities of scholarly retrieval systems, several studies have focused on the scholarly deep research [815], such as AcademicBrowse [932]. To facilitate section-level generation of related works, several large-scale datasets have been introduced, such as Cochrane [770], MSLR 2022 [787], MS2 [168], and OARelatedWork [173], OAG-Bench [891]. These datasets pair comprehensive “Related Work” sections with their corresponding full texts, providing valuable resources for this task. Moreover, systematic benchmarks for evaluating the quality of automatic scholarly survey generation have been developed. Examples include SciReviewGen [366], BigSurvey [491], SurveySum [205], SurveyBench [849], AutoSurvey [802], and SurveyX [462]. These benchmarks provide critical metrics for assessing the performance of automatic systems in generating academic surveys. For fine-grained manual annotation, SurveyEval [782] offers a hierarchical title tree that includes a vast number of reviews and citations. It is accompanied by hierarchical consistency and citation-chapter alignment metrics, which serve as essential tools for evaluating the distribution of synopsis generation and citation accuracy.

### v2

@@ 5.4.2 Pre-Experiment Estimation
predicts quantitative values or trends of experimental outcomes, such as estimating drug concentration effects, determining whether a compound affects cellular activity, and assessing protocol feasibility [153]. (1) Deep-Learning Methods: With the rise of deep learning, hierarchical prediction models have emerged [838]. Li et al. [439] incorporated physical equations to predict pharmacokinetic parameters, reducing data requirements and enhancing noise robustness. More recently, Li et al. [440] proposed a dual-matching framework, combining hierarchical molecular alignment with meta-learning, which showed significant improvements in drug feature estimation. (2) LLM-Augmented Methods: More recently, with the advent of LLM capabilities, Zhang et al. [909] demonstrate successful LLM-assisted evaluative prediction, a method that has been further extended in subsequent studies. Notably, Luo et al. [517] integrate BrainGPT into neuroscience literature retrieval, outperforming domain experts in evaluating experimental estimation. Wen et al. [822] develop a system combining fine-tuned GPT-4.1 with a paper retrieval agent, which outperformed 25 human experts in evaluating experimental predictions.
@@ 5.4.3 Experiment Management
5.4.4 Experiment Conduction
Experiment conduction refers to the application of AI techniques in executing and managing scientific experiments. This process is essential for automating workflows, ensuring experiments are carried out efficiently and accurately. The primary aim of experiment conduction is to reduce human involvement in the experimental process. It can be further divided into two categories:
@@ 5.4.4 Experiment Conduction
uses AI to streamline the design, training, and evaluation of ML models, reducing dependence on human expertise by covering the entire pipeline from preprocessing to hyperparameter optimization [905, 761, 256, 919, 576, 937]. Typically, Wang et al. [804] present a community-driven sandbox allowing agents to write code, browse the web, and coordinate through an event-stream API. For Kaggle challenges, Li et al. [459] propose an iterative, collaborative multi-agent system that incorporates debugging and unit testing across the competition pipeline. AIDE [176] utilizes a tree-search loop to generate, evaluate, and refine solutions, achieving a bronze medal in Kaggle competitions. At the research level, Li et al. [443] formalize a three-phase LLM agent workflow, idea generation, implementation, and execution, to automate experiments. Zhao et al. [929] and Liu et al. [483] present a multi-agent method for extracting model variables from scientific texts, significantly improving experimental reproduction accuracy. Cheng et al. [129] introduce a multi-agent framework for crafting novel LLMs by emulating standard research stages and leveraging scaling laws through a “Ladder of Scales”. Designs are proposed, adversarially reviewed, implemented, and evaluated across model sizes from 14 million to 350 million parameters.
@@
Recent advancements in the planning and reasoning capabilities of LLMs have led to their use in simulating experimental results [359, 889, 416, 888] and even direct conduct real-world experiments [123, 659, 203]. Real-world experimental simulation & conduction generally employ four strategies: (1) Self-Improvement: Models iteratively refine their performance based on feedback [313, 683, 916, 596, 883]. For example, Siddiqui et al. [688] enhance functional approximation through iterative knowledge application. Further refinement occurs through analytical insights [443, 737, 36] and hyperparameter tuning [568, 494, 901]. (2) Multi-Agent Interaction: Models simulate collaborative research teams by assigning roles such as experimenter, analyst, or critic [36, 234, 713, 307, 702]. For instance, MechAgent [567], Researchcodeagent [219] and The AI Scientist [509, 854] automate experiments through multi-agent collaboration, with LLMs acting as proxies in fields like computer science [509, 854], social science [552, 532], and physical science [906]. (3) External Tool Integration: Researchers enhance model capabilities by linking them to databases, APIs, and other tools during experiments [278, 621, 666]. For example, Boiko et al. [68] integrate internet search, code execution, and automation into a GPT-4 system. Studies like ChemCrow [520] and Crispr-GPT [316] support chemistry and gene editing experiments through massive specialized toolchains. (4) Specific Fine-Tuning: A growing body of work explores the fine-tuning of specific models to improve experimental simulations. For instance, Cui et al. [152] present a transformer-based model trained on large single-cell transcriptomic datasets, achieving state-of-the-art accuracy in cell-type annotation and in silico perturbation response [491].
@@ 6 AI for Academic Writing
AI for Academic writing involves the use of AI techniques to assist researchers or generate from scratch in drafting, editing, and formatting scientific manuscripts [373]. With the development of deeper interaction between human and LLMs, human and LLMs are quickly shaping each other’s better writing habits [230, 86, 943]. As shown in Figure 6, it contains two main categories: Semi-Automatic Academic Writing ($\S$ 6.1) and Full-Automatic Academic Writing ($\S$ 6.2).
@@ 6.1.2 Assistance During Manuscript Writing
serves as an effective means of conveying experimental data and analysis [55, 97]. Recent advancements in AI research on automatic scientific figure generation have led to significant breakthroughs [902, 295, 750, 638]. Rodriguez et al. [646] introduce the FigGen model, which maps textual descriptions to complex academic figures with high fidelity. Beyond direct image generation, several studies have explored programming scientific diagrams using Python [928, 921, 133, 553], SVG [648], or tikz [55, 54] for better figure quality. However, figures alone cannot fully convey results; captions are essential for ensuring that readers understand each figure [90, 509, 873]. To address this, Hsu et al. [302] develop SciCapenter, an interactive system that generates multiple caption candidates, scores them, and quality-checks each, assisting authors in selecting the most optimal phrasing. MLBCAP [377] utilize multiple LLMs to collaborate on chart-based title generation. Additionally, frameworks like AI Scientist [509, 854] can identify key data from experimental logs, generate figures with captions, and integrate them into authoring tools, advancing end-to-end AI-driven scientific visualization.
@@ 7 AI for Academic Peer Reviewing
Peer reviewing plays a crucial role in enhancing the quality of academic papers. However, it is often hindered by delays, time demands, and growing academic workloads [471, 386, 756, 956]. To address these challenges and improve dissertation quality, researchers are exploring the integration of AI into the review process [886, 492, 575, 401, 525]. As shown in Figure 7, it contains three main categories: Pre-Review ($\S$ 7.1), In-Review ($\S$ 7.2) and Post-Review ($\S$ 7.3).
@@ 7.1.1 Desk-Review
As manuscript submissions to academic journals increase, editorial offices face heavier workloads during the desk-review stage. To address this, major publishers have introduced AI-driven tools, such as automated keyword extraction, topic matching, and preliminary scoring, to enhance efficiency, shorten turnaround times, and reduce manual screening [178, 428, 198]. For example, Elsevier’s Evise and Editorial Manager (EM) systems use indexing and extract terms to route manuscripts to the appropriate subject areas and editorial teams [755]. Similarly, IEEE’s Manuscript Central (built on ScholarOne) combines metadata, author-provided keywords, and an academic-network reviewer discovery tool for more accurate matching [701]. Springer’s SNAPP system and Nature’s AI-assisted triage tool also demonstrate AI’s impact on desk-review workflows [562]. Additionally, D´ıaz et al. [171] develop AnnotateGPT, which generates annotations to help editors quickly assess a manuscript’s scope and quality, further speeding up the review process.
@@ 7.2 In-Review
Table 5: Comparison of expert and LLM review performance sourced from Shin et al. [682], where “GPT-4o-Finetuned” refers to GPT-4o finetuned with review data using the finetune-API. KL divergences are calculated from the average of four focus distributions (strength/target, weakness/target, strength/aspect, weakness/aspect) between expert and LLM reviews. F1 scores for overall performance, strength, and weakness are derived by comparing the (target, aspect) sets between expert and LLM reviews. Text similarity metrics are computed to assess the alignment between LLM and expert reviews. The bolded contents indicate the highest performance for each metric.
7.2.1 Peer-Review Generation
@@
integrates textual comments and numeric scores into a single review output that mirrors real-world peer review workflows [681, 417]. There are three main paradigms for optimizing unified peer-review generation: (1) Single-Agent Optimization: A straightforward approach is to optimize a single agent through deeper analysis [360, 950, 329]. Shin et al. [682] observe that, by comparing the focus distributions of LLMs and human experts, off-the-shelf LLMs tend to prioritize technical validity in paper reviews while underemphasizing novelty. To address this, Tyser et al. [765] enhance the review system with a suite of review documents to reduce risks of misuse, score inflation, overconfident assessments, and uneven distributions. Additionally, Zhu et al. [950], Zhang and Abernethy [907] incorporate deeper reasoning via reasoning LLMs to improve review quality. (2) Iterative Refinement Optimization: High-quality feedback is often ensured through hierarchical quality control and multi-round refinement loops [59]. Wu et al. [833] propose an LLM-driven pipeline with hierarchical verification, producing literature surveys that match human-authored reviews. Kirtani et al. [382] introduce standardized evaluation metrics and a self-refinement cycle to align LLM-generated reviews with human accuracy and analytical depth. (3) Multi-Agent Optimization: To further enhance feedback reliability, some studies adopt multi-agent frameworks [742, 323, 223, 572, 851]. D’Arcy et al. [159] divide manuscripts into modules for specialized agents, leading to higher-quality feedback than single-agent systems. CycleResearcher [825] and TreeReview [98] apply reinforcement learning to simulate iterative review rounds and structured agent interactions, enhancing collaboration. Furthermore, Taechoyotin and Acuna [738] propose multi-objective reinforcement learning to optimize unified peer review, while Taechoyotin et al. [739] extend multi-agent scientific reviews to multimodal scenarios.
7.2.2 Meta-Review Generation
@@ 7.3.1 Influence Analysis
Influence analysis seeks to predict the future scholarly impact of a paper, most commonly measured by citation count, by evaluating its intrinsic characteristics [697, 327]. Early approaches predominantly rely on external metadata or handcrafted features, such as author reputation and journal impact factor [174, 740, 954]. In contrast, recent methods leveraging LLMs offer the advantage of directly inferring a work’s innovativeness from its narrative. For instance, Zhao et al. [926] frame influence prediction as a regression task, fine-tuning an LLM on titles and abstracts to generate a time- and field-normalized impact score, effectively addressing the cold-start problem. Similarly, the HLM-Cite framework [277] adopts a two-stage approach: first, an embedding model retrieves a set of candidate citations from a large corpus, followed by a generative LLM that performs fine-grained reasoning and re-ranking to identify the most relevant references.
@@ 8 Application of AI for Research
As shown in Figure 8, it contains three main categories: AI for Natural Science Research ($\S$ 8.1), AI for Applied Science and Engineering Research ($\S$ 8.2) and AI for Social Science Research ($\S$ 8.3).
@@ 8.1.3 AI for Chemistry & Materials Research
seeks to identify optimal or novel material compositions in virtual or automated experimental setups while minimizing the number of required experiments [46, 633, 672]. Specifically, Chen et al. [102] introduce MEGNet, demonstrating that graph neural networks can achieve density functional theory-level accuracy for both molecular and crystalline properties. Li et al. [452] employ two-stage Bayesian optimization to screen 560 organic photocatalysts using only 2.4% of the experiment conditions, thereby obtaining significantly improved performance. Ekosso et al. [189] combine low-cost robotic platforms with high-throughput microscopy and Gaussian process models to map vesicle formation processes. More recently, Szymanski et al. [737] implement a robot-machine-learning platform that accelerate compound discovery and identify 41 novel inorganic materials within 17 days.
@@ 8.3.2 AI for Psychology Research
Research methodology focuses on the design, implementation, and validation of psychological experiments to ensure validity and reproducibility [760, 429, 899].
@@ 9.1.2 Table & Chart Scientific Comprehension
In the domain of reasoning based on charts and tables, a number of benchmarks have emerged to evaluate the ability of LLMs in both structural and logical comprehension. Early works, such as ChartQA [537], CharXiv [817], ChartX [842], MISSQA [931], and NovaChart [303], focus on assessing LLMs’ performance in answering questions related to charts, utilizing both synthetic and real-world data. On the other hand, benchmarks like SUC [721], TableBench [837], and ToRR [28] emphasize the evaluation of LLMs’ structural understanding, rather than content comprehension, across various tasks such as form interpretation, numerical reasoning, and textual analysis.
@@ 9.2 AI for Academic Survey
In the task of generating academic surveys, several representative public corpora stand out due to their distinct characteristics in terms of scale, domain, and structure. Table 7 presents a comprehensive overview of typical, mature AI systems and associated tools for academic survey collection and generation.
To evaluate the capabilities of scholarly retrieval systems, several studies have focused on the scholarly deep research [820], such as AcademicBrowse [942]. To facilitate section-level generation of related works, several large-scale datasets have been introduced, such as Cochrane [773], MSLR 2022 [791], MS2 [170], and OARelatedWork [175], OAG-Bench [898]. These datasets pair comprehensive “Related Work” sections with their corresponding full texts, providing valuable resources for this task. Moreover, systematic benchmarks for evaluating the quality of automatic scholarly survey generation have been developed. Examples include SciReviewGen [368], BigSurvey [493], SurveySum [208], SurveyBench [855], AutoSurvey [807], and SurveyX [464]. These benchmarks provide critical metrics for assessing the performance of automatic systems in generating academic surveys. For fine-grained manual annotation, SurveyEval [786] offers a hierarchical title tree that includes a vast number of reviews and citations. It is accompanied by hierarchical consistency and citation-chapter alignment metrics, which serve as essential tools for evaluating the distribution of synopsis generation and citation accuracy.